# Frontier AI’s Impact on the Cybersecurity Landscape

### v3

@@ 7.3 Build Secure AI Agents
We call for more effort in developing secure mechanisms for AI agents, whether through secure-by-design and secure-by-construction approaches or by creating specialized proactive and reactive defenses. This is essential to address the significant gap between the rapid emergence of AI agents and the limited exploration of their security. First, the security community has developed various protection mechanisms for traditional symbolic systems, such as strong authentication, access control, and secure protocol designs. However, with the rise of AI agents, there is currently no standardized secure design framework or principles tailored for such systems. This underscores the need for developing secure system design frameworks and principles for AI agents. These new frameworks and standards could adapt existing security enforcement mechanisms (e.g., privilege isolation), as well as create new mechanisms customized for the marginal risks unique to AI agents (e.g., tool permission controls and specific sandboxing). Having a more fine-grained rigorous isolation mechanism for AI systems also helps with safety and security, e.g., separating system prompts from user prompts helps mitigate prompt injections [38].
Second, we recommend building real-time monitoring and alert mechanisms for AI agents. To this end, we can extend existing AI attack detectors [8] to frontier AI, improve their efficiency, and continuously update them. In the long term, achieving provable security guarantees for AI agents is crucial. One possible direction is to follow a divide-and-conquer approach: decompose the verification of an AI agent into sub-goals for its AI and traditional symbolic components, applying or extending specific techniques to address each. Overall, this remains an open challenge requiring greater attention and collaborative efforts from multiple communities, including security, AI, and programming languages.
@@ 7.4 Recommendations for AI Development
Testing the cyberattack capabilities of new models before release and publishing detailed evaluation reports remains essential [113, 180]. As benchmarks evolve, AI developers should conduct more comprehensive testing beyond CTF competitions and collaborate with white-box hackers for real-world scenario testing. We can also explore implementing finer-grained privilege and access controls to ensure responsible use and avoid equivalence classes, enabling established defenders to access frontier AI more readily than attackers. However, implementation challenges remain, particularly in defining “established defenders.” Decisions about open model access require careful risk assessment [122, 190].
Safety and security demand continued and increased attention. We recommend extending current comprehensive red-teaming and scaled eval practices more broadly across AI development, particularly for AI agents. For blue-teaming, we advocate a best-effort approach incorporating complementary strategies: model safety alignment, guardrails for models and agents, and system-level protections. Transparency remains a key limitation of frontier AI [190]. Considering the challenges in explaining complex model decisions, transparency throughout the entire AI lifecycle—including development, training, deployment, and use—is crucial for informed policymaking. Implementing varied transparency levels for different entities (public, trusted third parties, government) could balance information sharing risks and benefits, though this approach may introduce implementation challenges and potential conflicts among stakeholders [190].
@@ 7.5 Mitigation of Human-Related Risks
We first call for better user and developer education about AI attacks and security practices. Human awareness remains critical in cybersecurity, with research showing that developers’ security knowledge gaps and user unawareness significantly increase vulnerability and damage [17, 19]. Studies demonstrate AI can enhance educational efficiency across domains [37, 195, 185], suggesting the potential of leveraging frontier AI for security education. AI-driven personalized learning systems can deliver more efficient security training tailored to individual knowledge levels [215, 138]. Intelligent nudge mechanisms represent another effective approach [97, 129], where AI systems detect potential phishing or scams and provide timely alerts, enhancing real-time threat recognition while minimizing user fatigue. However, education alone is insufficient, as users often remain the weakest link. AI-powered attacks affect humans on a massive scale, while defensive measures lag behind rapidly evolving threats. We urgently need defensive research addressing frontier AI-enabled sophisticated attacks rather than focusing solely on traditional scenarios.
@@
LLM-enhanced malicious bot activities, including misinformation campaigns and fraud schemes, exemplify this challenge. AI-based protective measures are essential against these advanced threats [74]. Such mechanisms can analyze behavioral patterns, communication styles, and interaction frequencies to distinguish legitimate users from sophisticated bots. Platforms can implement intelligent screening mechanisms to filter out malicious bots attempting social engineering or fraud on social media. Frontier AIs can create “honeypot” systems that attract and isolate these bots [188], engaging them in resource-draining interactions while learning their behaviors to develop better countermeasures. Given the rapid advancement of human-targeting attacks, game-theoretically based attack and defense analysis can provide valuable insights into this evolving landscape.
8 Discussion and Conclusion
@@
This paper presents an in-depth analysis of frontier AI’s current impacts on cybersecurity from three aspects (benchmarks, literature review, and expert survey), as well as projecting the future impact of AI and calls-to-action for steering AI to improve security. Even though current AI techniques have limitations, in this paper, we adopt an optimistic view: given the rapid evolution of models, AI holds significant potential to enhance the cybersecurity landscape. With this view, we provide recommendations on overcoming identified obstacles to promote AI in security. However, AI may fail, and there could be better alternative approaches. We leave the exploration of such alternatives for future work. This work highlights a few future directions. First, as a longitudinal study, we will continuously update our qualitative and quantitative analyses and expert survey to maintain an up-to-date understanding of AI’s evolving role in cybersecurity. Second, as AI is increasingly used in the physical world, future work can conduct a similar analysis for cyber-physical systems and broader hardware systems.
@@ Appendix A Performance of SOTA LLMs on Existing Attack and Defense Benchmarks
Table 5 and Table 6 summarize the performance of SOTA LLMs’ performance on popular attack and defense benchmarks discussed in the main text. In Table 5, all data are in percentage scale. The vulnerable code generation part reports the insecure code rates by instruction-based code generation. In the attack generation, we report the end-to-end attack success rate except for the RedCode-Gen which only involves code generation and we report the accuracy for it. For the SecCodePLT, we multiply the success rate of five steps to get the end-to-end success rate. For the CyBench and the NYU CTF benchmarks, we report the unguided solved rate and pass@1 rate separately. For the CyberBench benchmark, we report the average of the results in all sub-tasks. For the CyberMetric benchmark, we report the accuracy of the 2k questions. For the TACTL benchmark, we report the accuracy of the TACTL-183 benchmark. In Table 6, all data are in percentage scale. We report both autonomous (left) and assisted (right) success rates for the AutoPenBench. For the PrimeVul benchmark, we report the pair-wise correct prediction of the paired functions task with chain-of-thought. For the CRUXEval benchmark, we report the pass@1 rate on both input(left) and output(right) prediction results. For the SWE-bench-verified, we report the success rate. Note the agent architecture might be different.
@@
Table 5: Performance of SOTA LLMs on existing attack benchmarks. All data are in percentage scale with the value range between 0 and 100. ’-’ means the model is not tested on the corresponding benchmark.
  Models Vulnerable code generation Attack generation CTF Cyber knowledge
CyberSecEval-3 SecCodePLT CyberSecEval-3 SecCodePLT RedCode-Gen RedCode-Exec CyBench NYU CyberBench CyberMetric TACTL
gpt-3.5 - - - - 32.5 - - - 62.6 88.10 -
gpt-4 35 - 40 - 66.9 64.50 - 7.00 69.9 91.00 -
gpt-4o - 44 - 0.2 72.5 77.23 12.5 9.50 - 91.25 85.2
GPT-4.5-preview - - - - - - 17.5 - - - -
o1-preview - - - - - - 10.0 - - - -
o3-mini - - - - - - 22.5 - - - -
claude-3.5-sonnet - - - 0 - 67.63 17.5 13.50 - - -
Claude 3.7 Sonnet - - - - - - 20 - - - -
gemini-1.5-pro - - - - - - 7.5 - - - -
gemini-1.0-pro - - - - - - - - - 84.00 -
deepseek-r1 - - - - - - - - - - 91.8
deepseek-v3 - - - - - - - - - - 86.3
llama-3.1-405b 39 - 49 - - - 7.5 - - - 88.5
llama-3.3-70b - - - - - - - - - - 78.7
llama-3.1-70b 35 47 41 0 - 76.70 5.0 - - - -
llama-3.1-8b - - - - - 62.87 - - - - -
llama-3-8b - - - - - - - - - 73.05 -
llama-2-7b - - - - 20.7 - - - 50.6 - -
llama-2-70b - - - - - - - - - 71.60 -
@@
Table 6: Performance of SOTA LLMs on defense benchmarks. We report it only for the benchmarks for which results are publicly available. All data are in percentage scale with the value range between 0 and 100. ’-’ means the model is not tested on the benchmark. We report both autonomous (left) and assisted (right) success rates for the AutoPenBench.
  Models Pen. test Vul. detection PoC generation Patching
AutoPenBench PrimeVul CRUXEval SWE-bench-verified
gpt-3.5 - 6.21 49.1/63.3 0.4
gpt-4 - 12.94 74.8/81.9 22.4
gpt-4o 21.00/64.00 - - 38.8
o1 - - - 48.9
o3-mini - - - 49.3
claude-3.5-sonnet - - - 49.0
claude-3.7-sonnet - - - 70.3
deepseek-v3 - - - 42.0
deepseek-r1 - - - 49.2
code-llama-13b - - 39.1/39.3 -
code-llama-34b - - 50.4/46.0 -

### v4

@@ 8.3 Build Secure AI Agents
We call for more effort in developing secure mechanisms for AI agents, whether through secure-by-design and secure-by-construction approaches or by creating specialized proactive and reactive defenses. This is essential to address the significant gap between the rapid emergence of AI agents and the limited exploration of their security. First, the security community has developed various protection mechanisms for traditional symbolic systems, such as strong authentication, access control, and secure protocol designs. However, with the rise of AI agents, there is currently no standardized secure design framework or principles tailored for such systems. This underscores the need for developing secure system design frameworks and principles for AI agents. These new frameworks and standards could adapt existing security enforcement mechanisms (e.g., privilege isolation), as well as create new mechanisms customized for the marginal risks unique to AI agents (e.g., tool permission controls and specific sandboxing). Having a more fine-grained rigorous isolation mechanism for AI systems also helps with safety and security, e.g., separating system prompts from user prompts helps mitigate prompt injections [chen2024struq].
Second, we recommend building real-time monitoring and alert mechanisms for AI agents. To this end, we can extend existing AI attack detectors [aldahdooh2022adversarial] to frontier AI, improve their efficiency, and continuously update them. In the long term, achieving provable security guarantees for AI agents is crucial. One possible direction is to follow a divide-and-conquer approach: decompose the verification of an AI agent into sub-goals for its AI and traditional symbolic components, applying or extending specific techniques to address each. Overall, this remains an open challenge requiring greater attention and collaborative efforts from multiple communities, including security, AI, and programming languages.
@@ 8.4 Recommendations for AI Development
Testing the cyberattack capabilities of new models before release and publishing detailed evaluation reports remains essential [hurst2024gpt, phuong2024evaluating]. As benchmarks evolve, AI developers should conduct more comprehensive testing beyond CTF competitions and collaborate with white-box hackers for real-world scenario testing. We can also explore implementing finer-grained privilege and access controls to ensure responsible use and avoid equivalence classes, enabling established defenders to access frontier AI more readily than attackers. However, implementation challenges remain, particularly in defining “established defenders.” Decisions about open model access require careful risk assessment [kapoor2024societal, policy].
Safety and security demand continued and increased attention. We recommend extending current comprehensive red-teaming and scaled eval practices more broadly across AI development, particularly for AI agents. For blue-teaming, we advocate a best-effort approach incorporating complementary strategies: model safety alignment, guardrails for models and agents, and system-level protections. Transparency remains a key limitation of frontier AI [policy]. Considering the challenges in explaining complex model decisions, transparency throughout the entire AI lifecycle—including development, training, deployment, and use—is crucial for informed policymaking. Implementing varied transparency levels for different entities (public, trusted third parties, government) could balance information sharing risks and benefits, though this approach may introduce implementation challenges and potential conflicts among stakeholders [policy].
@@ 8.5 Mitigation of Human-Related Risks
We first call for better user and developer education about AI attacks and security practices. Human awareness remains critical in cybersecurity, with research showing that developers’ security knowledge gaps and user unawareness significantly increase vulnerability and damage [aslan2023comprehensive, bakhshi2017social]. Studies demonstrate AI can enhance educational efficiency across domains [chen2020artificial, santhi2024chat, rahman2023chatgpt], suggesting the potential of leveraging frontier AI for security education. AI-driven personalized learning systems can deliver more efficient security training tailored to individual knowledge levels [sreen2024leveraging, limo2023personalized]. Intelligent nudge mechanisms represent another effective approach [harrison2024behavioral, leal2024nudging], where AI systems detect potential phishing or scams and provide timely alerts, enhancing real-time threat recognition while minimizing user fatigue. However, education alone is insufficient, as users often remain the weakest link. AI-powered attacks affect humans on a massive scale, while defensive measures lag behind rapidly evolving threats. We urgently need defensive research addressing frontier AI-enabled sophisticated attacks rather than focusing solely on traditional scenarios.
@@
LLM-enhanced malicious bot activities, including misinformation campaigns and fraud schemes, exemplify this challenge. AI-based protective measures are essential against these advanced threats [ferrara2023social]. Such mechanisms can analyze behavioral patterns, communication styles, and interaction frequencies to distinguish legitimate users from sophisticated bots. Platforms can implement intelligent screening mechanisms to filter out malicious bots attempting social engineering or fraud on social media. Frontier AIs can create “honeypot” systems that attract and isolate these bots [trapbot], engaging them in resource-draining interactions while learning their behaviors to develop better countermeasures. Given the rapid advancement of human-targeting attacks, game-theoretically based attack and defense analysis can provide valuable insights into this evolving landscape.
9 Concluding Remarks
@@
This paper presents an in-depth analysis of frontier AI’s current impacts on cybersecurity from four aspects: benchmarks, literature review, empirical evaluation, and expert survey. Even though current AI techniques have limitations, in this paper, we adopt an optimistic view: given the rapid evolution of models, AI holds significant potential to enhance the cybersecurity landscape. With this view, we provide recommendations on overcoming identified obstacles to promote AI in security. However, AI may fail, and there could be better alternative approaches. We leave the exploration of such alternatives for future work. This work also points to several promising directions for future research. First, as part of a longitudinal effort, we plan to continuously update our analyses to capture the evolving role of AI in cybersecurity. Second, as AI becomes more deeply integrated into physical systems, future studies could extend this analysis to domains such as cyber-physical systems and broader hardware infrastructures.
@@ Appendix A Performance of SOTA LLMs on Existing Attack and Defense Benchmarks
Table 6 and Table 7 summarize the performance of SOTA LLMs’ performance on popular attack and defense benchmarks discussed in the main text. In Table 6, all data are in percentage scale. The vulnerable code generation part reports the insecure code rates by instruction-based code generation. In the attack generation, we report the end-to-end attack success rate except for the RedCode-Gen which only involves code generation and we report the accuracy for it. For the SeCodePLT, we multiply the success rate of five steps to get the end-to-end success rate. For the CyBench and the NYU CTF benchmarks, we report the unguided solved rate and pass@1 rate separately. For the CyberBench benchmark, we report the average of the results in all sub-tasks. For the CyberMetric benchmark, we report the accuracy of the 2k questions. For the TACTL benchmark, we report the accuracy of the TACTL-183 benchmark. Overall the SOTA agents achieve better performance on the earlier attack steps than later steps. This is aligned with our qualitative analysis and reflect the difficult of later steps. The domain-specific tools are gdb, hydra, pwntools, metasploit, ghidra, tshark for existing attack agents.
In Table 7, all data are in percentage scale. We report both autonomous (left) and assisted (right) success rates for the AutoPenBench. For the PrimeVul benchmark, we report the pair-wise correct prediction of the paired functions task with chain-of-thought. For the SeCodePLT benchmark, we report the Python w/o policy rate for vulnerability detection and report the pass@1 rate for the Patching. For the CRUXEval benchmark, we report the pass@1 rate on both input(left) and output(right) prediction results. For the CyberGym benchmark, we report the pass@1 rate. For the SWE-bench-verified, we report the success rate. Overall, the agents perform poorly on vulnerability detection, PoC generation, and sub-optimal on patching. Compared to attack agents, defense agents are equipped with fewer domain-specific tools. Program analysis tools, which are supposed to be useful, are not equiped in existing agents.
@@
Table 6: Performance of SOTA LLMs on existing attack benchmarks. All data are in percentage scale with the value range between 0 and 100. N/A means model only without agent scaffold. $\uparrow$ indicates that the higher the value, the better the performance, and vice versa. The domain-specific tools in CyBench agent are pwntools, metasploit, ghidra, tshark.
  Vulnerable code generation Attack generation
CyberSecEval-3 $\downarrow$ SeCodePLT $\downarrow$ CyberSecEval-3 $\uparrow$ SeCodePLT $\uparrow$ RedCode-Gen $\uparrow$ RedCode-Exec $\uparrow$
Models gpt-4 turbo claude-sonnet-3.7 llama-3.1-405b gpt-4o deepseekcoder-6.7b deepseekcoder-6.7b
Tools N/A N/A N/A domain-specific (kali) N/A general
Performance 34 11 49 0.2 79.4 80.23
  CTF Cyber knowledge
CyBench $\uparrow$ NYU $\uparrow$ CyberBench $\uparrow$ CyberMetric $\uparrow$ TACTL $\uparrow$
Models claude-sonnet-4.5 + kali-based env claude-sonnet-3.5 + CRAKEN gpt-4 gpt-4o deepseek-r1
Tools domain-specific domain-specific (gdb) N/A N/A N/A
Performance 55 22 69.6 91.25 91.8
@@
Table 7: Performance of SOTA LLMs on defense benchmarks. We report it only for the benchmarks for which results are publicly available. All data are in percentage scale with the value range between 0 and 100. ’N/A’ means no agent scaffold is used. We report both autonomous (left) and assisted (right) success rates for the AutoPenBench, and report pass@1 rate on both input (left) and output (right) for the CRUXEval. The domain-specific tools used is the kali in the CyberGym agent.
  Vul. detection PoC generation Patching
PrimeVul SeCodePLT CRUXEval CyberGym SeCodePLT
Models gpt-4 o4-mini gpt-4-turbo + cot claude-sonnet-4.5 + kali-based env claude-sonnet-3.7
Tools N/A N/A N/A domain-specific (kali) N/A
Performance 12.94 65.2 75.7/82.0 28.9 63.9